# Quasi-convex Splittings of Acylindrical Graphs of Locally finite-height Groups

### v1

@@ 2.1. Finite-height and Quasi-Convex Splittings
Let $G$ be a group with bounded finite subgroup, and assume that $G$ acts on a tree $T$ with finitely many orbits of edges. If the action of $G$ on $T$ is a finite-height splitting then it is also an acylindrical splitting.
As stated above, the conditions in the statement of Lemma 2.11 are not overly restrictive. Indeed, these conditions can always be assumed to hold for hyperbolic groups — a hyperbolic group is finitely generated so any action on a tree admits a unique minimal subtree with finitely many orbits of edges by Lemma 2.1, and a hyperbolic group will always have bounded finite subgroups by [Bogopolskii95, Theorem], for example. Motivated by Lemma 2.11 we make the following definition.
@@
Let $G$ be a group with bounded finite subgroups, and assume that $G$ admits a finite-height splitting with finitely many orbits of edges. This splitting will be acylindrical by Lemma 2.11, so we say that this splitting is non-elementary if it is non-elementary in the sense of Theorem 2.6
The concept of a finite-height splitting is strictly stronger than that of an acylindrical one in the setting of groups with bounded finite subgroups acting on trees with finitely many orbits of edges. Consider the following example.
@@
Then by [Cohen23, Corollary 3.12], the splitting given by $\langle a,b,c\rangle*_{\langle b,c\rangle}\langle b,c,d\rangle$ is non-elementary acylindrical, but the edge group does not have finite-height in $G$. Indeed, the conjugate of $\langle b,c\rangle$ by any power of $a$ will contain the copy of $F_{2}$ generated by $b$ and $c$, so the height of $\langle b,c\rangle$ in $G$ will be infinite.
In this example, finite-height fails “locally” — the edge group is not finite-height in the vertex groups into which it includes. It turns out that this is the only way in which an acylindrical splitting can fail to be finite-height, as shown by the following lemma.
@@
Let $G$ be a group acting $(k,C)$-acylindrically on some tree $T$ with finitely many orbits of edges. Then this action is a finite-height splitting if and only if each edge group has finite-height in both of its neighbours, i.e. for all $e\in E(\Gamma)$, $h_{G_{i(e)}}(d_{e,0}(G_{e}))<\infty$ and $h_{G_{t(e)}}(d_{e,1}(G_{e}))<\infty$.
@@
Firstly, for the if direction assume that the action of $G$ on $T$ is a finite-height splitting. Then the stabiliser $G_{e}$ of each edge $e\in E(T)$ must have finite-height in $G$, so in turn must have finite-height in the vertex stabilisers of the endpoints of $e$ by definition.
@@
We separate into two cases. First, if the graph theoretical diameter of $S$ is at least $k$ then $|\operatorname{PStab}_{G}(S)|\leq C$ by the acylindricity assumption, so in particular is finite. For the second case, assume the graph theoretical diameter of $S$ is less than $k$. As the diameter of $S$ is less than $k$, $S$ must have some vertex $v$ of degree at least $n$ as a tree of diameter less than $k$ and maximum vertex degree $n-1$ has less than $(n-1)^{k}$ vertices, and so less than $(n-1)^{k}-1<n^{k}$ edges. Thus by construction of $n$ we have that $\operatorname{PStab}_{G}(S)=\bigcap_{i=1}^{n^{k}}g_{i}G_{e}g_{i}^{-1}$ is finite.
Consequently in both cases $\bigcap_{i=1}^{n^{k}}g_{i}G_{e}g_{i}^{-1}$ is finite, and so this is always finite and $G_{e}$ has height at most $n^{k}$ as claimed. It follows that each edge stabiliser of the action of $G$ on $T$ has has height at most $n^{k}$, and therefore has finite-height as required. ∎
@@
Many examples of quasi-convex subgroups of hyperbolic groups exist in the literature. For example, since every bounded set is quasi-isometric to the point, all finite subgroups of any hyperbolic group are quasi-convex, and it is a well known result (see Bowditch [Bowditch98cutpoints, Section 1], for example) that every two-ended subgroup of a hyperbolic group is quasi-convex. Further important examples of quasi-convex subgroup arise as geometrically finite subgroups of fundamental groups of closed and orientable hyperbolic $3$-manifolds, see [Hruska08, Corollary 1.3].
We say that the action of a hyperbolic group $G$ on a tree $T$ is a quasi-convex splitting if the $G$-stabiliser of every edge $e\in E(T)$ is quasi-convex in $G$.
@@
Many important examples of quasi-convex splittings are well known. Perhaps the most important example is the JSJ-decomposition of a hyperbolic group, which is the “maximal” splittings of a $1$-ended hyperbolic group along two ended subgroups [Bowditch98cutpoints].
@@
    [BridsonHaefliger, Proposition 3.7] The subgroups $H_{1}$ and $H_{2}$ are hyperbolic, so in particular they are finitely generated.
@@
    [BridsonHaefliger, Proposition 3.9] The intersection $H_{1}\cap H_{2}$ is a quasi-convex subgroup of $G$.
The link to finite-height splittings, and therefore acylindrical splittings, comes from the following influential result of Gitik, Mj, Rips and Sageev.
@@
[GitikMahanRipsSageev98, Main Theorem] Let $G$ be a hyperbolic group, and $H\leq G$ a quasi-convex subgroup of $G$. Then $H$ has finite-height in $G$.
@@
By Theorem 2.19 the action of $G$ on $T$ is finite-height, and so recalling that minimal actions of hyperbolic groups on trees will always satisfy the other conditions of Lemma 2.11 we obtain that the action of $G$ on $T$ is acylindrical. Finally, the action will be non-elementary by Theorem 2.6 and the fact that the action is non-trivial and $G$ is not virtually cyclic. ∎
2.2. The Dunwoody Resolution
@@
Let $(n_{i})_{i=1}^{\infty}$ be the sequence of natural numbers defined inductively by $n_{1}=1$, $n_{i+1}=456+8n_{i}^{2}$. Define $x_{i}=ab^{n_{i}}ab^{n_{i}+1}\cdots ab^{n_{i+1}-1}$, and $X_{i}\subset\langle a,b\rangle$ by $X_{i}=\{x_{1},...,x_{i}\}$. By choice of the $n_{i}$s each $X_{i}$ has the property that $\langle a,b\mid X_{i}\rangle$ is a $C^{\prime}(1/6)$ presentation, and so by [LOGAN2019116, Theorem F] we have that for all $i$, $\langle X_{i}\rangle\leq\langle a,b\rangle$ is a malnormal free subgroup of $\langle a,b\rangle$ of rank $i$ whose basis is $X_{i}$. It follows that $H=\langle\bigcup_{i\in\mathbb{N}}X_{i}\rangle\leq\langle a,b\rangle$ is an infinitely generated malnormal subgroup of $\langle a,b\rangle$.
We may then decompose $F_{3}$ as the amalgam $\langle a,b\rangle*_{H}(H*\langle c\rangle)$. The subgroup $H$ is malnormal in $\langle a,b\rangle$, which is itself malnormal in $G$, so $H$ is malnormal in $G$ and thus has height $2$ in $G$. This decomposition therefore corresponds to a finite-height splitting of a hyperbolic group over an infinitely generated edge group. In particular, by Lemma 2.18(1), $H$ is not quasi-convex in $F_{3}$, and so there exists a finite-height splitting of a hyperbolic group which is not a quasi-convex splitting.
@@
We may always construct actions with finitely generated edge stabilisers from splittings of finitely presented groups using a construction known as the Dunwoody resolution.
@@
[DunwoodyResolution, Lemma 2.2] If $G$ is a finitely presented group acting by simplicial isometry on a tree $T$ without inversion, then there exists a tree $T^{\prime}$ such that $G$ acts on $T^{\prime}$ by simplicial isometry and without inversion with finitely generated edge stabilisers, and such that there exists a $G$-equivariant graph epimorphism
@@
In general, if we have two $G$-trees $T^{\prime}$ and $T$ related by such an epimorphism we say that $T^{\prime}$ dominates $T$. Dunwoody’s original proof is constructive, and works by pulling back the midpoints of edges of the tree $T$ onto the Cayley complex associated to some finite presentation of $G$ to give nicely embedded 1-complexes called tracks, which we can cut along to give our new splitting. The exact construction will not be of use in this paper, but can be found in [DunwoodyResolution].
@@ 3. Acylindrical Actions with Well Behaved Stabilisers
We wish to use the Dunwoody resolution to construct quasi-convex (or finite-height with finitely generated edge stabilisers) splittings of hyperbolic groups from acylindrical ones, and as such we are particularly interested in the conditions under which it will preserve acylindricity.
@@ 3.1. Local Properties of Groups
The main result of this paper is to prove that if we are given some degree of control over the stabilisers of our action then the Dunwoody Resolution will preserve acylindricity. In particular, we will seek to control the local properties of such stabilisers.
@@
    Finitely generated free groups, surface groups, and more generally hyperbolic limit groups [Wilton2006HallsTF, Corollary 3.12] and many small cancellation groups [MccamondWiseLocalQC, MccamondWiseLocalQC2] are locally quasi-convex. It follows that splittings of such groups over finitely generated subgroups will always be quasi-convex splittings, and so if such a group admits any non-trivial action on a tree, then it will admit a quasi-convex splitting arising from the Dunwoody resolution.
@@
    We say that a group $G$ is locally finite-height if every finitely generated subgroup of $G$ has finite-height in $G$. By Theorem 2.19 a locally quasi-convex group will always be locally finite-height, but locally finite-height groups form a much larger class as they are not required to be hyperbolic, or even finitely generated. For example, the free group on countably many generators is locally finite-height owing to its inclusion into $F_{2}$, which is locally quasi-convex.
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite-height then $G$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers that dominates the action of $G$ on $T$.
@@
    If there exists an edge stabiliser that is finite-height in $G$, and this edge stabiliser is locally finite-height, then $G$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers. Furthermore, if every edge stabiliser satisfies this condition then we may assume that the new action dominates the action of $G$ on $T$.
@@
We begin with a proof of point (a). By Theorem 2.22 there exists some tree $T^{\prime}$ on which $G$ acts with finitely generated edge stabilisers (and therefore finitely generated vertex stabilisers by Lemma 2.2) such that $T^{\prime}$ dominates $T$ along the $G$-equivarient epimorphism $\phi:T^{\prime}\rightarrow T$. By Lemma 2.1 the action of $G$ on $T^{\prime}$ has a unique minimal subtree $T^{\prime}_{m}$ with finitely many orbits of edges, and the image of the restriction of $\phi$ to $T^{\prime}_{m}$ is a fixed subtree of $T$. This restriction must therefore be surjective by the fact that the action of $G$ on $T$ was minimal, and thus the action of $G$ on $T^{\prime}_{m}$ dominates the action of $G$ on $T$. We will therefore assume that $T^{\prime}=T^{\prime}_{m}$ for the remainder of this proof, and we will show that $T^{\prime}$ corresponds to a non-elementary finite-height splitting with finitely generated edge stabilisers.
@@
Now let $u\in U_{v}$. The group $\operatorname{PStab}_{G_{v}}(u)$ is a finitely generated subgroup of $G_{v}$, so must have finite-height in $G_{v}$ by the assumption that $G_{v}$ is locally finite-height. Define $n_{u}$ to be the height of $\operatorname{PStab}_{G_{v}}(u)$ in $G_{v}$. Pick a set of representatives $S\subset U_{v}$ of the $G_{v}$-orbits in $U_{v}$, and define
@@
The action of $G$ on $T^{\prime}$ is therefore $(k^{\prime},C_{G})$-acylindrical, and thus it only remains to show that the action of $G$ on $T^{\prime}$ is a finite-height splitting. For all $v\in V(T^{\prime})$ the stabiliser $\operatorname{PStab}_{G}(v)$ is a subgroup of the stabiliser $\operatorname{PStab}_{G}(\phi(v))$, and so in particular is locally finite-height. Thus, as each edge stabiliser is finitely generated, this action will satisfy the conditions of Lemma 2.11 and will be a finite-height splitting as required. It only remains to check that the action of $G$ on $T^{\prime}$ is non-elementary. The group $G$ is acylindrically arboreal by assumption, so $G$ is not virtually cyclic and the action of $G$ on $T^{\prime}$ cannot be lineal by Theorem 2.6. The action of $G$ on $T^{\prime}$ can also have no fixed points, as it dominates the action of $G$ on $T$ with has no fixed points, and so the action of $G$ on $T^{\prime}$ will be non-elementary by Theorem 2.6 as required.
@@
We now move onto a proof of part b of this theorem. The proof of this is similar to part a, but we include it for completeness and to make clear the small differences. We begin by modifying $T$ by collapsing the quotient graph of groups to consider only the splitting over the edges $e\in E(\Gamma)$ with the properties required in the hypothesis of part b, and call this new graph of groups $(\Gamma_{c},\mathfrak{G}_{c})$. This will still correspond to a minimal non-elementary acylindrical splitting by [Cohen23, Lemma 2.20]. Construct the tree $T^{\prime}$ as before, which we will again assume to be minimal with a finite number of orbits of edges. Therefore $T^{\prime}$ will dominate our original splitting if no edge of $\Gamma$ was collapsed, as required. Again as before, fix some $e\in\Gamma_{c}$ and consider the set of edges $S_{e}=\phi^{-1}(e)\subseteq E(T^{\prime})$, which we may equip with a metric given by considering the distance between their midpoints in $T^{\prime}$. By equivariance, $G_{e}=\operatorname{PStab}_{G}(e)$ acts on $S_{e}$ by isometry and the action has properties analogous to properties (1) and (2) above.
Now let $f\in S_{e}$. The group $\operatorname{PStab}_{G_{e}}(f)$ is a finitely generated subgroup of $G_{e}$, so must have finite-height in $G_{e}$ by the assumption that $G_{e}$ is locally finite-height. Define $n_{f}$ to be the height of $\operatorname{PStab}_{G_{e}}(f)$ in $G_{e}$. Pick a set of representatives $R\subset S_{e}$ of the $G_{e}$-orbits in $S_{e}$, and define
$k_{e}:=1+\sum_{f\in R}n_{f},$
which is finite by point (2), and define $k_{edge}=\max_{e\in e\Gamma_{c}}(k_{e})$. Define
$k^{\prime}:=\left(1+\sum_{e\in E(\Gamma_{c})}h_{G}(G_{e})\right)\cdot k_{edge}.$
Finally, let $P\in T^{\prime}$ be a set of at least $k^{\prime}$ edges in $E(T^{\prime})$ (as opposed to the path chosen in the proof of part a). We then have two cases, the first of which occurs when $\phi(P)$ contains at least $1+\sum_{e\in E(\Gamma_{c})}h_{G}(G_{e})$ edges. Then the stabiliser $\operatorname{PStab}_{G}(P)$ is contained in the stabiliser $\operatorname{PStab}_{G}(\phi(P)$, which in turn is contained in the intersection of at least $h_{G}(G_{e})$ essentially disjoint $G$-conjugates of $G_{e}$ for some $e\in E(\Gamma_{c})$ by the pigeonhole principle. It follows by the definition of $h_{G}(G_{e})$ that $\operatorname{PStab}_{G}(P)$ finite, and so contains at most $C_{G}$ elements.
The second case occurs when $\phi(P)$ contains fewer than $1+\sum_{e\in E(\Gamma_{c})}h_{G}(G_{e})$ edges, so at least one edge $e\in\phi(P)$ has at least $k_{edge}$ preimages in $P$ by the pigeonhole principle. My construction $k_{edge}\geq k_{e}$, so again by the pigeonhole principle there exists some orbit $G_{e}\cdot f$ of edges in $\phi^{-1}(e)$ such that $|\phi^{-1}(e)\cap P\cap(G_{e}\cdot f)|\geq n_{f}$. It follows that, since each edge in $\phi^{-1}(e)\cap P\cap(G_{e}\cdot f)$ is pointwise stabilised (with respect to the $G$-action on $T^{\prime}$) by a $G_{e}$-conjugate of $\operatorname{PStab}_{G_{e}}(f)$, the $G$-stabiliser of this set must be finite, and therefore so must $\operatorname{PStab}_{G}(P)$.
We thus have that every set of edges of $T^{\prime}$ of size at least $k^{\prime}$ must have a finite stabiliser in $G$, and so each edge stabiliser of the action of $G$ on $T^{\prime}$ must have height at most $k^{\prime}$ in $G$. In particular, this action is a finite-height splitting for $G$ as required, and will be non-elementary as required as in the proof of part (a). ∎
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite-height, then $H$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers that dominates the induced action of $H$ on $T_{H}$.
@@
    If there exists an edge $e\in E(T_{H})$ whose $G$-stabiliser (with respect to the action of $G$ on $T$) has finite-height in $G$, and this edge stabiliser is locally finite-height, then $H$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers. Furthermore, if every edge $e\in E(T_{H})$ satisfies this condition then we may assume that the new action dominates the induced action of $H$ on $T_{H}$.

### v2

@@ 2. Preliminaries - Acylindrical Hyperbolicity and Acylindrical Arboreality
Let $G$ be a group with bounded finite subgroup, and assume that $G$ acts on a tree $T$ with finitely many orbits of edges. If the action of $G$ on $T$ is a finite height splitting then it is also an acylindrical splitting.
As stated above, the conditions in the statement of Lemma 2.11 are not overly restrictive. Indeed, these conditions can always be assumed to hold for hyperbolic groups — a hyperbolic group is finitely generated so any action on a tree admits a unique minimal subtree with finitely many orbits of edges by Lemma 2.1, and a hyperbolic group will always have bounded finite subgroups by [undefd, Theorem], for example. Motivated by Lemma 2.11 we make the following definition.
@@
Let $G$ be a group with bounded finite subgroups, and assume that $G$ admits a finite height splitting with finitely many orbits of edges. This splitting will be acylindrical by Lemma 2.11, so we say that this splitting is non-elementary if it is non-elementary in the sense of Theorem 2.6
The concept of a finite height splitting is strictly stronger than that of an acylindrical one in the setting of groups with bounded finite subgroups acting on trees with finitely many orbits of edges. Consider the following example.
@@
Then by [undefi, Corollary 3.12], the splitting given by $\langle a,b,c\rangle*_{\langle b,c\rangle}\langle b,c,d\rangle$ is non-elementary acylindrical, but the edge group does not have finite height in $G$. Indeed, the conjugate of $\langle b,c\rangle$ by any power of $a$ will contain the span of $b$, so the height of $\langle b,c\rangle$ in $G$ will be infinite.
In this example, finite height fails “locally” — the edge group is not finite height in the vertex groups into which it includes. It turns out that this is the only way in which an acylindrical splitting can fail to be finite height, as shown by the following lemma.
@@
Let $G$ be a group acting $(k,C)$-acylindrically on some tree $T$ with finitely many orbits of edges. Then this action is a finite height splitting if and only if each edge group has finite height in both of its neighbours, i.e. for all $e\in E(\Gamma)$, $h_{G_{i(e)}}(d_{e,0}(G_{e}))<\infty$ and $h_{G_{t(e)}}(d_{e,1}(G_{e}))<\infty$.
@@
Firstly, for the if direction assume that the action of $G$ on $T$ is a finite height splitting. Then the stabiliser $G_{e}$ of each edge $e\in E(T)$ must have finite height in $G$, so in turn must have finite height in the vertex stabilisers of the endpoints of $e$ by definition.
@@
We separate into two cases. First, if the graph theoretical diameter of $S$ is at least $k$ then $|\operatorname{PStab}_{G}(S)|\leq C$ by the acylindricity assumption, so in particular is finite. For the second case, assume the graph theoretical diameter of $S$ is less than $k$. As the diameter of $S$ is less than $k$, $S$ must have some vertex $v$ of degree at least $n$ as a tree of diameter less than $k$ and maximum vertex degree n1n-1 has less than $(n-1)^{k}$ vertices, and so less than $(n-1)^{k}-1<n^{k}$ edges. Thus by construction of $n$ we have that $\operatorname{PStab}_{G}(S)=\bigcap_{i=1}^{n^{k}}g_{i}G_{e}g_{i}^{-1}$ is finite.
Consequently in both cases $\bigcap_{i=1}^{n^{k}}g_{i}G_{e}g_{i}^{-1}$ is finite, and so this is always finite and $G_{e}$ has height at most $n^{k}$ as claimed. It follows that each edge stabiliser of the action of $G$ on $T$ has has height at most $n^{k}$, and therefore has finite height as required. ∎
@@
Many examples of quasi-convex subgroups of hyperbolic groups exist in the literature. For example, since every bounded set is quasi-isometric to the point, all finite subgroups of any hyperbolic group are quasi-convex, and it is a well known result (see Bowditch [undefe, Section 1], for example) that every two-ended subgroup of a hyperbolic group is quasi-convex. Further important examples of quasi-convex subgroup arise as geometrically finite subgroups of fundamental groups of closed and orientable hyperbolic $3$-manifolds, see [undefp, Corollary 1.3].
We say that the action of a hyperbolic group $G$ on a tree $T$ is a quasi-convex splitting if the $G$-stabiliser of every edge $e\in E(T)$ is quasi-convex in $G$. In such a splitting the vertex stabilisers are also strongly controlled. We have the following result of Bowditch.
Lemma 2.17.
[undefe, Proposition 1.2] Let $G$ be a hyperbolic group acting on a tree $T$ with quasi-convex edge stabilisers and finitely many orbits of edges. Then each vertex stabiliser of this action is also quasi-convex in $G$.
@@
Many important examples of quasi-convex splittings are well known. Perhaps the most important example is the JSJ-decomposition of a hyperbolic group, which is the “maximal” splittings of a $1$-ended hyperbolic group along two ended subgroups [undefe].
@@
    [undefg, Proposition 3.7] The subgroups $H_{1}$ and $H_{2}$ are hyperbolic, so in particular they are finitely generated.
@@
    [undefg, Proposition 3.9] The intersection $H_{1}\cap H_{2}$ is a quasi-convex subgroup of $G$.
The link to finite height splittings, and therefore acylindrical splittings, comes from the following influential result of Gitik, Mj, Rips and Sageev.
@@
[undefm, Main Theorem] Let $G$ be a hyperbolic group, and $H\leq G$ a quasi-convex subgroup of $G$. Then $H$ has finite height in $G$.
@@
By Theorem 2.20 the action of $G$ on $T$ is finite height, and so recalling that minimal actions of hyperbolic groups on trees will always satisfy the other conditions of Lemma 2.11 we obtain that the action of $G$ on $T$ is acylindrical. Finally, the action will be non-elementary by Theorem 2.6 and the fact that the action is non-trivial and $G$ is not virtually cyclic. ∎
2.2. The Dunwoody–Sageev Resolution
@@
Let $(n_{i})_{i=1}^{\infty}$ be the sequence of natural numbers defined inductively by $n_{1}=1$, $n_{i+1}=456+8n_{i}^{2}$. Define $x_{i}=ab^{n_{i}}ab^{n_{i}+1}\cdots ab^{n_{i+1}-1}$, and $X_{i}\subset\langle a,b\rangle$ by $X_{i}=\{x_{1},...,x_{i}\}$. By choice of the $n_{i}$s each $X_{i}$ has the property that $\langle a,b\mid X_{i}\rangle$ is a $C^{\prime}(1/6)$ presentation, and so by [undefv, Theorem F] we have that for all $i$, $\langle X_{i}\rangle\leq\langle a,b\rangle$ is a malnormal free subgroup of $\langle a,b\rangle$ of rank $i$ whose basis is $X_{i}$. It follows that $H=\langle\bigcup_{i\in\mathbb{N}}X_{i}\rangle\leq\langle a,b\rangle$ is an infinitely generated malnormal subgroup of $\langle a,b\rangle$.
We may then decompose $F_{3}$ as the amalgam $\langle a,b\rangle*_{H}(H*\langle c\rangle)$. The subgroup $H$ is malnormal in $\langle a,b\rangle$, which is itself malnormal in $G$, so $H$ is malnormal in $G$ and thus has height $2$ in $G$. This decomposition therefore corresponds to a finite height splitting of a hyperbolic group over an infinitely generated edge group. In particular, by Lemma 2.19(1), $H$ is not quasi-convex in $F_{3}$, and so there exists a finite height splitting of a hyperbolic group which is not a quasi-convex splitting.
@@
We may always construct actions with finitely generated edge stabilisers from splittings of finitely presented groups using a construction known as the Dunwoody–Sageev resolution.
@@
[undefk, Lemma 2.2] If $G$ is a finitely presented group acting by simplicial isometry on a tree $T$ without inversion, then there exists a tree $T^{\prime}$ such that $G$ acts on $T^{\prime}$ by simplicial isometry and without inversion with finitely generated edge stabilisers, and such that there exists a $G$-equivariant graph epimorphism
@@
In general, if we have two $G$-trees $T^{\prime}$ and $T$ related by such an epimorphism we say that $T^{\prime}$ dominates $T$. Dunwoody and Sageev’s original proof is constructive, and works by pulling back the midpoints of edges of the tree $T$ onto the Cayley complex associated to some finite presentation of $G$ to give nicely embedded 1-complexes called tracks, which we can cut along to give our new splitting. The exact construction will not be of use in this paper, but can be found in [undefk].
@@ 3. Acylindrical Actions with Well Behaved Stabilisers
We wish to use the Dunwoody–Sageev resolution to construct quasi-convex (or finite height with finitely generated edge stabilisers) splittings of hyperbolic groups from acylindrical ones, and as such we are particularly interested in the conditions under which it will preserve acylindricity.
@@ 3.1. Local Properties of Groups
The main result of this paper is to prove that if we are given some degree of control over the stabilisers of our action then the Dunwoody–Sageev Resolution will preserve acylindricity. In particular, we will seek to control the local properties of such stabilisers.
@@
    Finitely generated free groups, surface groups, and more generally hyperbolic limit groups [undefaj, Corollary 3.12] and many small cancellation groups [undefz, undefaa] are locally quasi-convex. It follows that splittings of such groups over finitely generated subgroups will always be quasi-convex splittings, and so if such a group admits any non-trivial action on a tree, then it will admit a quasi-convex splitting arising from the Dunwoody–Sageev resolution.
@@
    We say that a group $G$ is locally finite height if every finitely generated subgroup of $G$ has finite height in $G$. By Theorem 2.20 a locally quasi-convex group will always be locally finite height, but locally finite height groups form a much larger class as they are not required to be hyperbolic, or even finitely generated. For example, the free group on countably many generators is locally finite height owing to its inclusion into $F_{2}$, which is locally quasi-convex.
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite height then $G$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers that dominates the action of $G$ on $T$.
@@
    If there exists an edge stabiliser that is finite height in $G$, and this edge stabiliser is locally finite height, then $G$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers. Furthermore, if every edge stabiliser satisfies this condition then we may assume that the new action dominates the action of $G$ on $T$.
@@
We begin with a proof of point (a). By Theorem 2.23 there exists some tree $T^{\prime}$ on which $G$ acts with finitely generated edge stabilisers (and therefore finitely generated vertex stabilisers by Lemma 2.2) such that $T^{\prime}$ dominates $T$ along the $G$-equivariant epimorphism $\phi:T^{\prime}\rightarrow T$. By Lemma 2.1 the action of $G$ on $T^{\prime}$ has a unique minimal subtree $T^{\prime}_{m}$ with finitely many orbits of edges, and the image of the restriction of $\phi$ to $T^{\prime}_{m}$ is a fixed subtree of $T$. This restriction must therefore be surjective by the fact that the action of $G$ on $T$ was minimal, and thus the action of $G$ on $T^{\prime}_{m}$ dominates the action of $G$ on $T$. We will therefore assume that $T^{\prime}=T^{\prime}_{m}$ for the remainder of this proof, and we will show that $T^{\prime}$ corresponds to a non-elementary finite height splitting with finitely generated edge stabilisers.
@@
Now let $u\in U_{v}$. The group $\operatorname{PStab}_{G_{v}}(u)$ is a finitely generated subgroup of $G_{v}$, so must have finite height in $G_{v}$ by the assumption that $G_{v}$ is locally finite height. Define $n_{u}$ to be the height of $\operatorname{PStab}_{G_{v}}(u)$ in $G_{v}$. Pick a set of representatives $S\subset U_{v}$ of the $G_{v}$-orbits in $U_{v}$, and define
@@
The action of $G$ on $T^{\prime}$ is therefore $(k^{\prime},C_{G})$-acylindrical, and thus it only remains to show that the action of $G$ on $T^{\prime}$ is a finite height splitting. For all $v\in V(T^{\prime})$ the stabiliser $\operatorname{PStab}_{G}(v)$ is a subgroup of the stabiliser $\operatorname{PStab}_{G}(\phi(v))$, and so in particular is locally finite height. Thus, as each edge stabiliser is finitely generated it must have finite height in both of its neighbours, and the action of $G$ on $T^{\prime}$ will be a finite height splitting by Lemma 2.14 as required. It only remains to check that the action of $G$ on $T^{\prime}$ is non-elementary. The group $G$ is acylindrically arboreal by assumption, so $G$ is not virtually cyclic and the action of $G$ on $T^{\prime}$ cannot be lineal by Theorem 2.6. The action of $G$ on $T^{\prime}$ can also have no fixed points, as it dominates the action of $G$ on $T$ with has no fixed points, and so the action of $G$ on $T^{\prime}$ will be non-elementary by Theorem 2.6 as required.
@@
We now move onto a proof of part b of this theorem. We begin by modifying $T$ by collapsing the quotient graph of groups to consider only the splitting over the edges $e\in E(\Gamma)$ with the properties required in the hypothesis of part b, and call this new graph of groups $(\Gamma_{c},\mathfrak{G}_{c})$. This will still correspond to a minimal non-elementary acylindrical splitting by [undefi, Lemma 2.20], with associated Bass-Serre tree $T_{c}$. Construct the tree $T^{\prime}$ as before using Theorem 2.23 applied to $T_{c}$, which we will again assume to be minimal with a finite number of orbits of edges. Therefore $T^{\prime}$ will dominate our original splitting if no edge of $\Gamma$ was collapsed, as required. Again as before, fix some $e\in E(T_{c})$, and consider the set of edges $S_{e}=\phi^{-1}(e)\subseteq E(T^{\prime})$, which we may equip with a metric given by considering the distance between their midpoints in $T^{\prime}$. By equivariance, $G_{e}=\operatorname{PStab}_{G}(e)$ acts on $S_{e}$ by isometry and the action has the properties analogous to property (1) above.
Now let $f\in E(T^{\prime})$. By property (1), the stabiliser $\operatorname{PStab}_{G}(f)$ agrees with the stabiliser $\operatorname{PStab}_{G_{\phi(f)}}(f)$, which is therefore a finitely generated subgroup of $G_{\phi(f)}$ and thus has finite height in $G_{\phi(f)}$. In turn, $G_{\phi(f)}$ has finite height in $G$ by hypothesis, and so by transitivity of the finite height property the stabiliser $\operatorname{PStab}_{G}(f)$ must have finite height in $G$. The action of $G$ on $T^{\prime}$ is therefore a finite height splitting, and will therefore be acylindrical by Lemma 2.11. Finally, the action of $G$ on $T^{\prime}$ will be non-elementary as required as in the proof of part a.
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite height, then $H$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers that dominates the induced action of $H$ on $T_{H}$.
@@
    If there exists an edge $e\in E(T_{H})$ whose $G$-stabiliser (with respect to the action of $G$ on $T$) has finite height in $G$, and this edge stabiliser is locally finite height, then $H$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers. Furthermore, if every edge $e\in E(T_{H})$ satisfies this condition then we may assume that the new action dominates the induced action of $H$ on $T_{H}$.